# Pumping and Steady Streaming driven by Two-Frequency Oscillations of a Cylinder

### v1

@@ A.2. Solution Structure of $p_{n}$
$\begin{array}[]{ll}\displaystyle\frac{\partial p_{n}}{\partial r}=&\Delta u_{r,n}-\frac{u_{r,n}}{r^{2}}-\frac{2}{r^{2}}\frac{\partial u_{\theta,n}}{\partial\theta}\\[10.0pt] &-\text{Re}\displaystyle\frac{\partial u_{r,n}}{\partial t}-\text{Re}\sum_{i+j=n}\left(u_{r,i}\frac{\partial u_{r,j}}{\partial r}+\frac{u_{\theta,i}}{r}\frac{\partial u_{r,j}}{\partial\theta}-\frac{u_{\theta,i}u_{\theta,j}}{r}\right),\end{array}\\ &&\begin{array}[]{ll}\displaystyle\frac{1}{r}\frac{\partial p_{n}}{\partial\theta}=&\Delta u_{\theta,n}-\frac{u_{\theta,n}}{r^{2}}+\frac{2}{r^{2}}\frac{\partial u_{r,n}}{\partial\theta}-\text{Re}\frac{\partial u_{\theta,n}}{\partial t}\\[10.0pt] &-\text{Re}\displaystyle\sum_{i+j=n}\left(u_{r,i}\frac{\partial u_{\theta,j}}{\partial r}+\frac{u_{\theta,i}}{r}\frac{\partial u_{\theta,j}}{\partial\theta}+\frac{u_{r,i}u_{\theta,j}}{r}\right).\end{array}$ (60)
@@
$\frac{\partial p_{n}}{\partial r}&=&L_{r,n}+\text{Re}\sum_{i+j=n}NL_{r,i,j},\\ \frac{1}{r}\frac{\partial p_{n}}{\partial\theta}&=&L_{\theta,n}+\text{Re}\sum_{i+j=n}NL_{\theta,i,j},$ (61)
@@
$L_{r,n}$ $=\Delta u_{r,n}-\frac{u_{r,n}}{r^{2}}-\frac{2}{r^{2}}\frac{\partial u_{\theta,n}}{\partial\theta}-\text{Re}\frac{\partial u_{r,n}}{\partial t},$ (62a)
$L_{\theta,n}$ $=\Delta u_{\theta,n}-\frac{u_{\theta,n}}{r^{2}}+\frac{2}{r^{2}}\frac{\partial u_{r,n}}{\partial\theta}-\text{Re}\frac{\partial u_{\theta,n}}{\partial t}$ (62b)
@@
$p_{n}=\begin{cases}\sum_{l=0}^{n/2}G_{2l}^{n}(r,t)\cos(2l\theta)&\textbf{for n even}\\ \\ \sum_{l=1}^{(n-1)/2}G_{2l+1}^{n}(r,t)\cos\left((2l+1)\theta\right)&\textbf{for n odd}.\end{cases}$ (63)
Proof: We will examine the case when $n$ is odd, as showing the case when $n$ is even entails the same procedure. To prove the result, we show that the $r$-component of Equation (60) can be expressed as a linear combination of $\{\cos(\theta),\cos(3\theta),\cdots,\cos(n\theta)\}$, and the $\theta$-component can be expressed as a linear combination of $\{\sin(\theta),\sin(3\theta),\cdots,\sin(n\theta)\}$. We show this is true first for the linear terms, and then for the nonlinear terms.
@@
$M_{m}^{n}(r,t)=-\frac{4m}{r^{3}}\left(-\frac{r^{2}}{4}\frac{\partial^{2}F_{m}^{n}}{\partial r^{2}}-\frac{r}{4}\frac{\partial F_{m}^{n}}{\partial r}+\frac{m^{2}}{4}F_{m}^{n}+\frac{\text{Re}\;r^{2}}{4}\frac{\partial F_{m}^{n}}{\partial t}\right).$
@@
$S_{s}^{n}(r,t)=-\frac{\partial^{3}}{\partial r^{3}}F_{s}^{n}-\frac{1}{r}\frac{\partial^{2}F_{s}^{n}}{\partial r^{2}}+\left(\frac{s}{r}\right)^{2}\frac{\partial F_{s}^{n}}{\partial r}+\frac{1}{r^{2}}\frac{\partial F_{s}^{n}}{\partial r}-\frac{2s^{2}}{r^{3}}F_{s}^{n}+\text{Re}\frac{\partial^{2}F_{s}^{n}}{\partial t\partial r}.$
@@
Finally, from (64) and (68), $\partial p_{n}/\partial r$ is composed of cosine functions with odd arguments, and similarly, from (65) and (70), $\partial p_{n}/\partial\theta$ is composed of sine functions with even arguments. Integrating with respect to both $r$ and $\theta$, the solution structure of $p_{n}$ is
$p_{n}=\sum_{l=1}^{(n-1)/2}G_{2l+1}^{n}(r,t)\cos((2l+1)\theta).$
@@ Appendix B Establishing Necessary Conditions for Pumping
In Section 4.4, we showed that for frequency ratio 2, the third order terms proportional to $\sin(\theta)$ in the steady, stream function (and $\cos(\theta)$ in the steady pressure) produce a net force. To extend the argument from Section 4.4 to general frequency ratios, we derive necessary conditions for the presence of a steady term proportional to $\sin(\theta)$ in the stream function or proportional to $\cos(\theta)$ in the pressure. Lemmas A.1 and A.2 state that $\sin(\theta)$ and $\cos(\theta)$ terms can only appear when $n$ is odd. Therefore, the necessary condition for pumping is equivalent to finding order $n$ so that $n$ is odd, and $\psi_{n}$ has a steady term.
@@ B.1. Frequencies at each order
$\Delta p=-\text{Re}\;\boldsymbol{\nabla}\boldsymbol{\cdot}\left(\boldsymbol{u}\cdot\boldsymbol{\nabla}\boldsymbol{u}\right).$ (76)
@@
$\Delta p_{n}=-\text{Re}\sum_{i+j=n}\boldsymbol{\nabla}\boldsymbol{\cdot}\left(\boldsymbol{u}_{i}\boldsymbol{\cdot}\boldsymbol{\nabla}\boldsymbol{u}_{j}\right).$ (77)
@@ B.2. Existence of Pumping
Without a loss of generality choose $a$ odd and $b$ even. From (79), steady terms occur at odd order $n$ for $c$ odd. The smallest odd order $n$ is obtained by choosing $c=1$ and $\xi_{2}=\xi_{3}=0$, which is order $n=a+b$.
@@ C.1. Immersed Boundary Method
$\boldsymbol{f}(\boldsymbol{x},t)=S\boldsymbol{F}=\int_{\text{structure}}\boldsymbol{F}(s,t)\delta(\boldsymbol{x}-\boldsymbol{X}(\boldsymbol{s},t))\;ds,$ (80)
@@
In our simulations the motion of the structure is prescribed, and the force density on the structure, $\boldsymbol{F}(s,t)$, is determined implicitly by requiring the fluid velocity match the prescribed velocity of the boundary:
@@
$\rho\left(\frac{\partial\boldsymbol{u}}{\partial t}+\boldsymbol{u}\boldsymbol{\cdot}\boldsymbol{\nabla}\boldsymbol{u}\right)$ $=-\boldsymbol{\nabla}p+\mu\Delta\boldsymbol{u}+S\boldsymbol{F},$ (82)
$\boldsymbol{\nabla}\boldsymbol{\cdot}\boldsymbol{u}$ $=0,$ (83)
@@ C.2. Discretization
For the temporal discretization we use a second-order IMEX scheme, named SBDF in Ascher et al. (1995), in which the nonlinear terms are treated explilty in time, and the terms for the viscous force and structure force are treated implicitly in time with BDF2. The discretized system is
$\rho\left(\frac{3\boldsymbol{u}^{n+1}-4\boldsymbol{u}^{n}+\boldsymbol{u}^{n-1}}{2\Delta t}+2\boldsymbol{u}^{n}\boldsymbol{\cdot}G\boldsymbol{u}^{n}-\boldsymbol{u}^{n-1}\boldsymbol{\cdot}G\boldsymbol{u}^{n-1}\right)$ $=-Gp^{n+1}+\mu L\boldsymbol{u}^{n+1}+S\boldsymbol{F}^{n+1},$ (86)
@@
$g=\rho\left(-2\boldsymbol{u}^{n}\boldsymbol{\cdot}G\boldsymbol{u}^{n}+\boldsymbol{u}^{n-1}\boldsymbol{\cdot}G\boldsymbol{u}^{n-1}-\frac{-4\boldsymbol{u}^{n}+\boldsymbol{u}^{n-1}}{2\Delta t}\right).$
@@
The grid resolution for all simulations was $\Delta x=8/256$, so that there were $256$ grid points in the vertical direction, or equivalently $64$ points along the cylinder diameter. The time step was $\Delta{t}=0.0005$. The tolerance for the conjugate gradient method used to solve (90) was $0.001$.

### v2

@@ A.2. Solution Structure of $p_{n}$
$\frac{\partial p_{n}}{\partial r}&=&\Delta u_{r,n}-\frac{u_{r,n}}{r^{2}}-\frac{2}{r^{2}}\frac{\partial u_{\theta,n}}{\partial\theta}-Re\frac{\partial u_{r,n}}{\partial t}-Re\sum_{i+j=n}\left(u_{r,i}\frac{\partial u_{r,j}}{\partial r}+\frac{u_{\theta,i}}{r}\frac{\partial u_{r,j}}{\partial\theta}-\frac{u_{\theta,i}u_{\theta,j}}{r}\right),\\ \frac{1}{r}\frac{\partial p_{n}}{\partial\theta}&=&\Delta u_{\theta,n}-\frac{u_{\theta,n}}{r^{2}}+\frac{2}{r^{2}}\frac{\partial u_{r,n}}{\partial\theta}-Re\frac{\partial u_{\theta,n}}{\partial t}-Re\sum_{i+j=n}\left(u_{r,i}\frac{\partial u_{\theta,j}}{\partial r}+\frac{u_{\theta,i}}{r}\frac{\partial u_{\theta,j}}{\partial\theta}+\frac{u_{r,i}u_{\theta,j}}{r}\right).$ (62)
@@
$\frac{\partial p_{n}}{\partial r}&=&L_{r,n}+Re\sum_{i+j=n}NL_{r,i,j},\\ \frac{1}{r}\frac{\partial p_{n}}{\partial\theta}&=&L_{\theta,n}+Re\sum_{i+j=n}NL_{\theta,i,j},$ (63)
@@
$L_{r,n}$ $=\Delta u_{r,n}-\frac{u_{r,n}}{r^{2}}-\frac{2}{r^{2}}\frac{\partial u_{\theta,n}}{\partial\theta}-Re\frac{\partial u_{r,n}}{\partial t},$ (64a)
$L_{\theta,n}$ $=\Delta u_{\theta,n}-\frac{u_{\theta,n}}{r^{2}}+\frac{2}{r^{2}}\frac{\partial u_{r,n}}{\partial\theta}-Re\frac{\partial u_{\theta,n}}{\partial t}$ (64b)
@@
$p_{n}=\begin{cases}\sum_{l=0}^{n/2}G_{2l}^{n}(r,t)\cos(2l\theta)&\textbf{for n even}\\ \\ \sum_{l=0}^{(n-1)/2}G_{2l+1}^{n}(r,t)\cos\left((2l+1)\theta\right)&\textbf{for n odd}.\end{cases}$ (65)
Proof: We will examine the case when $n$ is odd, as showing the case when $n$ is even entails the same procedure. To prove the result, we show that $r$-component of (62) can be expressed as a linear combination of $\{\cos(\theta),\cos(3\theta),\cdots,\cos(n\theta)\}$ and the $\theta$-component can be expressed as a linear combination of $\{\sin(\theta),\sin(3\theta),\cdots,\sin(n\theta)\}$. We show this is true first for the linear terms, and then for the nonlinear terms.
@@
$M_{m}^{n}(r,t)=-\frac{4m}{r^{3}}\left(-\frac{r^{2}}{4}\frac{\partial^{2}F_{m}^{n}}{\partial r^{2}}-\frac{r}{4}\frac{\partial F_{m}^{n}}{\partial r}+\frac{m^{2}}{4}F_{m}^{n}+\frac{Re\;r^{2}}{4}\frac{\partial F_{m}^{n}}{\partial t}\right).$
@@
$S_{s}^{n}(r,t)=-\frac{\partial^{3}}{\partial r^{3}}F_{s}^{n}-\frac{1}{r}\frac{\partial^{2}F_{s}^{n}}{\partial r^{2}}+\left(\frac{s}{r}\right)^{2}\frac{\partial F_{s}^{n}}{\partial r}+\frac{1}{r^{2}}\frac{\partial F_{s}^{n}}{\partial r}-\frac{2s^{2}}{r^{3}}F_{s}^{n}+Re\frac{\partial^{2}F_{s}^{n}}{\partial t\partial r}.$
@@
Finally, from (66) and (70), $\partial p_{n}/\partial r$ is composed of cosine functions with odd arguments, and similarly, from (67) and (72), $\partial p_{n}/\partial\theta$ is composed of sine functions with odd arguments. Integrating with respect to both $r$ and $\theta$, the solution structure of $p_{n}$ is
$p_{n}=\sum_{l=0}^{(n-1)/2}G_{2l+1}^{n}(r,t)\cos((2l+1)\theta).$
@@ Appendix B Establishing Necessary Conditions for Pumping
In Section 4.4, we showed that for frequency ratio $\alpha=2$, the third order terms proportional to $\sin(\theta)$ in the steady, stream function (and $\cos(\theta)$ in the steady pressure) produce a net force. To extend the argument from Section 4.4 to general frequency ratios, we derive necessary conditions for the presence of a steady term proportional to $\sin(\theta)$ in the stream function or proportional to $\cos(\theta)$ in the pressure. Lemmas A.1 and A.2 state that $\sin(\theta)$ and $\cos(\theta)$ terms can only appear when $n$ is odd. Therefore, the necessary condition for pumping is equivalent to finding order $n$ so that $n$ is odd, and $\psi_{n}$ has a steady term.
@@ B.1. Frequencies at each order
$\Delta p=-Re\;\boldsymbol{\nabla}\boldsymbol{\cdot}\left(\boldsymbol{u}\boldsymbol{\cdot}\boldsymbol{\nabla}\boldsymbol{u}\right).$ (78)
@@
$\Delta p_{n}=-Re\sum_{i+j=n}\boldsymbol{\nabla}\boldsymbol{\cdot}\left(\boldsymbol{u_{i}}\boldsymbol{\cdot}\boldsymbol{\nabla}\boldsymbol{u_{j}}\right).$ (79)
@@ B.2. Existence of Pumping
Without a loss of generality, choose $a$ odd and $b$ even. From (81), steady terms occur at odd order $n$ for $c$ odd. The smallest odd order $n$ is obtained by choosing $c=1$ and $\xi_{2}=\xi_{3}=0$, which is order $n=a+b$.
@@ C.1. Immersed Boundary Method
$\boldsymbol{f}(\boldsymbol{x},t)=S\boldsymbol{F}=\int_{\text{structure}}\boldsymbol{F}(s,t)\delta(\boldsymbol{x}-\boldsymbol{X}(s,t))\;ds,$ (82)
@@
In our simulations, the motion of the structure is prescribed, and the force density on the structure, $\boldsymbol{F}(s,t)$, is determined implicitly by requiring the fluid velocity match the prescribed velocity of the boundary:
@@
$\rho\left(\frac{\partial\boldsymbol{u}}{\partial t}+\boldsymbol{u}\cdot\nabla\boldsymbol{u}\right)$ $=-\nabla p+\mu\Delta\boldsymbol{u}+S\boldsymbol{F},$ (84)
$\nabla\cdot\boldsymbol{u}$ $=0,$ (85)
@@ C.2. Discretization
For the temporal discretization, we use a second-order IMEX scheme, named SBDF in Ascher et al. (1995), in which the nonlinear terms are treated explilty in time, and the terms for the viscous force and structure force are treated implicitly in time with BDF2. The discretized system is
$\rho\left(\frac{3\boldsymbol{u}^{n+1}-4\boldsymbol{u}^{n}+\boldsymbol{u}^{n-1}}{2\Delta t}+2\boldsymbol{u}^{n}\cdot G\boldsymbol{u}^{n}-\boldsymbol{u}^{n-1}\cdot G\boldsymbol{u}^{n-1}\right)$ $=-Gp^{n+1}+\mu L\boldsymbol{u}^{n+1}+S\boldsymbol{F}^{n+1},$ (88)
@@
$g=\rho\left(-2\boldsymbol{u}^{n}\cdot G\boldsymbol{u}^{n}+\boldsymbol{u}^{n-1}\cdot G\boldsymbol{u}^{n-1}-\frac{-4\boldsymbol{u}^{n}+\boldsymbol{u}^{n-1}}{2\Delta t}\right).$
@@
The grid resolution for channel simulations was $\Delta x=8/256$, so that there were $256$ grid points in the vertical direction, or equivalently $64$ points along the cylinder diameter. The time step was $\Delta{t}=5.00\times 10^{-4}$. The tolerance for the conjugate gradient method used to solve (92) was $1\times 10^{-3}$.
$\Delta x$ $\Delta t$ Flux Difference Ratio $\log_{2}(\text{Ratio})$
8/256 $1.00\times 10^{-2}$ $9.6044\times 10^{-2}$
8/512 $5.00\times 10^{-3}$ $1.0130\times 10^{-1}$ $5.2573\times 10^{-3}$
8/1024 $2.50\times 10^{-3}$ $1.0279\times 10^{-1}$ $1.4937\times 10^{-3}$ 3.5197 1.8155
8/2048 $1.25\times 10^{-3}$ $1.0324\times 10^{-1}$ $4.4149\times 10^{-4}$ 3.3833 1.7584
Table 1. Refinement study for the flux with simultaneous refinement of space and time for the flow around an oscillating cylinder in a square 8 by 8 domain with periodic boundary conditions at $\text{Re}=10$, amplitude $\epsilon=0.5$, and time T = 100.
For the simulations presented in Figure 1, the grid resolution was $\Delta x=8/512$ and the time step was $\Delta{t}=5.00\times 10^{-3}$. Table 1 shows the results of a mesh refinement study on this problem which shows that the time-averaged flux converges at approximately second-order. Using the finest mesh $(\Delta x=8/2048,\Delta t=1.25\times 10^{-3})$ as the solution, we estimate the error in the flux for $\Delta x=8/512$ and $\Delta t=5.00\times 10^{-3}$ is approximately $1.9\%$.
$\Delta t$ Flux Difference Ratio $\log_{2}(\text{Ratio})$
$1.000\times 10^{-2}$ $9.6044\times 10^{-2}$
$5.000\times 10^{-3}$ $1.0086\times 10^{-1}$ $4.8185\times 10^{-3}$
$2.500\times 10^{-3}$ $1.0175\times 10^{-1}$ $8.8700\times 10^{-4}$ 5.4324 2.4416
$1.250\times 10^{-3}$ $1.0191\times 10^{-1}$ $1.6291\times 10^{-4}$ 5.4447 2.4448
$6.250\times 10^{-4}$ $1.0194\times 10^{-1}$ $3.2425\times 10^{-5}$ 5.0243 2.3289
$3.125\times 10^{-4}$ $1.0195\times 10^{-1}$ $7.1196\times 10^{-6}$ 4.5543 2.1872
Table 2. Refinement study in time only for the flux with the space step fixed at $\Delta x=8/256$ for the flow around an oscillating cylinder in a square 8 by 8 domain with periodic boundary conditions at $\text{Re}=10$, amplitude $\epsilon=0.5$, and time T = 100.
In the channel simulations, we used a coarser spatial resolution, $\Delta x=8/256$, but a smaller time step $\Delta t=5.00\times 10^{-4}$. Fixing the grid resolution as $\Delta x=8/256$ and refining the time step, we find that the flux converges approximately second-order in $\Delta t$ (Table 2). For the smallest time step, the relative changes in flux are on the scale of $10^{-4}$, and thus the spatial error is dominant. Comparing with the fluxes from the space-time refinement study in Table 1, we estimate the error in our numerical simulations to be on the scale of $1\--2\%$.